OpenStudy (anonymous):

y + 3 + 2y - 4 = 6 ?

OpenStudy (anonymous):

collect like terms

OpenStudy (mimi_x3):

3y -1=6 3y = 7 y = 7/3

OpenStudy (saifoo.khan):

What was your answer?

OpenStudy (mimi_x3):

woops my bad 3y = -7 y = -7/3

OpenStudy (anonymous):

I am stuckbecause it didnt have a number first

OpenStudy (mimi_x3):

you collect the like terms and numbers xD

OpenStudy (anonymous):

Im confused

OpenStudy (anonymous):

y+2y=6+4-3 which is 3y=7 which implies y=7/3